OpenStudy (anonymous):

True or False. The following lines are parallel. 2x + 2y = 4 and y = -x + 5

OpenStudy (texaschic101):

parallel lines will have the same slope but different y intercepts. 2x + 2y = 4 -- divide by 2 x + y = 2 y = -x + 2 slope is -1 and y intercept is 2 y = -x + 5 slope = -1 and y intercept is 5 so what do you think the answer is now ?

OpenStudy (anonymous):

so it would be true thanks